# Learning-Based Difficulty Calibration for Enhanced Membership Inference Attacks

### v1

@@ 4.3 Effect of data augmentation
Data augmentation is a technique that can be used to increase the size of a dataset by applying different transformations to existing training data. This technique is often used to improve the generalization and robustness of models. By exposing the model to a wider range of data variations, it can learn to handle different input scenarios, resulting in better performance. In this experiment, we applied data augmentation during the training of the target model and the shadow models (if any), and then evaluated its effect on the MIAs by measuring its AUC and TPR at $1\%$ FPR values. Note that we used horizontal flipping for the shadow (target) models training, and random cropping and rotation for the target model training. This is important because attackers may not have access to the data augmentation techniques used in the target model in real-life situations. The results of the experiment are shown in Table IV and Table V, where we only present the results for the CIFAR-10 and CIFAR-100 datasets due to space limitations.
TABLE IV: The impact of data augmentation on AUC
@@
w/o aug w aug w/o aug w aug
Salem et al. [30] 0.703 0.612 0.697 0.625
Yeom et al. [39] 0.694 0.601 0.889 0.802
Watson et al. [37] 0.711 0.667 0.772 0.709
Ours 0.754 0.719 0.918 0.869
TABLE V: The impact of data augmentation
on TPR at $1\%$ FPR
Attack method CIFAR-10 CIFAR-100
w/o aug w aug w/o aug w aug
Salem et al. [30] 1.6% 0.5% 2.3% 0.6%
Yeom et al. [39] 0.0% 0.0% 1.0% 0.0%
Watson et al. [37] 3.5% 1.1% 5.9% 2.7%
Ours 6.4% 3.6% 16.4% 12.4%
@@
4.4 Comparison with LiRA
The likelihood ratio attack (LiRA) [3] is a state-of-the-art MIA that can achieve a high TPR at low FPR regions. To determine membership, LiRA calculates the likelihood ratio, which represents the ratio of the likelihood of a data point being in the target model’s training dataset to the likelihood of it being from a different, unknown dataset. The likelihood ratio is computed based on the output probabilities of $N$ shadow models. The membership is determined based on the likelihood ratio. However, online LiRA attack is computationally intensive, because, for every new (batch) of target instance(s), LiRA needs to train multiple new shadow and reference models, with and without these instances. Note that LDC-MIA only utilizes two auxiliary models: the shadow target model and the reference model in the experiments. The same neural network classifier in LDC-MIA, trained with data provided by the shadow target model and reference model, can be re-used and attack new target instances without any need for additional model training.
TABLE VI: TPR at Low FPR regions on CIFAR-10.
Attack method ResNet-18 ResNet-34
0.1% 1% 0.1% 1%
LiRA(N=2) 1.5% 4.6% 1.2% 4.7%
LiRA(N=4) 1.5% 5.6% 1.7% 5.3%
LiRA(N=6) 1.7% 5.8% 2.0% 5.9%
LiRA(N=8) 1.9% 5.7% 2.0% 6.0%
Ours 2.6% 10.7% 2.1% 8.8%
In the comparison experiment, we select ResNet-18 and ResNet-34 trained on CIFAR-10 as the target models. The TPR at low FPR regions are shown in Table VI, both LDC-MIA and LiRA.
4.5 Cost-Performance Trade-off
In real attack scenarios, the cost associated with an attack is also a crucial factor. Thus, in this section, we compare the trade-off between cost and performance among the MIAs that we have evaluated. We use the TPR at $0.1\%$ FPR as the metric to evaluate the performance of MIAs. On the other hand, We use the number of auxiliary models required to launch the attack as the proxy metrics for the cost of the attack, this metric approximates both the computational cost and the data cost.
Table VII includes the TPR results obtained from attacking the CIFAR-10 dataset with ResNet-18 as the target model, and the number of shadow and reference models required for each attack method. It is important to note that the LiRA method needs to train 8 (or 256, depending on the setting) shadow models for every target instance. Thus, the total cost is not a constant, but rather linear to the number of target instances. We also note that our method LDC-MIA needs to train a neural network-based MIA classifier in addition to one shadow model and one reference model. This model uses a shallow MLP architecture. It poses a similar or less cost than training a shadow model in all of the datasets we experiment, and significantly less in the case where the shadow model is a DNN-based image model.
TABLE VII: The Cost-Performance Trade-off of MIAs.
Attack Shadow Reference TPR
Method Models Models at $0.1\%$FPR
LiRA[3](N=8) 8 - 1.9%
LiRA[3](N=256) 256 - 8.1%
Watson et al. [37](N=1) - 1 0.9%
Watson et al. [37](N=10) - 10 1.4%
Salem et al. [30] 1 - 0.2%
Yoem et al. [39] - - 0.0%
Ours 1 1 2.6%
The table shows that LiRA can achieve the highest TPR if 256 shadow models are trained for each target sample. However, our proposed attack method reduces the cost of attacking significantly. Meanwhile, it still manages to achieve a higher TPR than all other existing MIAs. The results indicate that LDC-MIA strikes an outstanding balance between performance and cost.
@@ 5.1 Differential Privacy
In order to evaluate the robustness of our proposed attack, we utilize the concept of differential privacy during the training of the target model. This technique adds noise or randomization to the data, which helps protect individual privacy in datasets. It can be useful in limiting the effectiveness of many existing MIAs [37, 3, 39]. We use DP-SGD [1], one of the state-of-the-art mechanisms, for training the target model in our experiments. DP-SGD adds carefully calibrated noise to the gradients computed during each iteration. The amount of noise added depends on the sensitivity of the gradients and the desired privacy budget $\varepsilon$. While a smaller $\varepsilon$ provides stronger privacy guarantees, it can also result in noisier updates. The other two important parameters in DP-SGD are the clipping bound $C$ and the noise multiplier $\sigma$.
@@
From the figure and the table, it is evident that as the desired privacy budget $\varepsilon$ increases, both AUC and TPR decrease. However, in practice, there is a trade-off between privacy ($\varepsilon$) and utility (the accuracy of the trained model). When higher privacy is required, adding more noise can significantly affect the model’s utility. This is evident from the model accuracy column in Table VIII. In practice, to preserve model accuracy, reasonable values of $\varepsilon$, such as 100 or 1000, are more likely to be used. It is observed that AUC and TPR are not significantly reduced in these cases. Moreover, even with small $\varepsilon$ values, the proposed attack can still achieve high precision values ($>90\%$) at low recall, as seen from Figure 9(b).
@@ 5.2 Overfitting Level of the Target Model
Previous studies [30, 32] have demonstrated that the performance of MIAs is closely related to the overfitting level of the target model. Overfitting happens when the model learns the training data too well, including noise and specific patterns that are unique to the training set. There are several factors that can affect the overfitting level of a model, such as dataset size and quality, training duration, model complexity, and regularization. To vary the overfitting level of the target model in our experiments, we adjust the training dataset size between 6500 and 12500. Typically, increasing the size of the training dataset helps reduce overfitting, while decreasing it has the opposite effect. A larger and more diverse dataset allows the model to learn a wider range of patterns and variations present in the data. Meanwhile, we keep the size of the training dataset of the reference model $\mathcal{D}^{train}_{ref}$ and that of the shadow target model $\mathcal{D}^{train}_{shadow}$ fixed. We then measure the impact of the overfitting level by evaluating AUC and TPR of the proposed attack. The results are shown in Table IX.
@@
In Table IX, we can observe that the overfitting level of the target model increases as the size of its training dataset decreases. This can be seen from the gap between the training accuracy and test accuracy of the target model. The larger the gap between these two, the more the model is overfitting. Furthermore, we can also see that the AUC and TPR at $1\%$ FPR improve slightly as the overfitting level of the target model increases.
@@ 5.3 Training Dataset Sizes for the Shadow Target and the Reference Models
An important factor in our proposed attack is the size of the datasets used to train the shadow target and reference models. As discussed in Section 5.2, the training dataset sizes affect the performance of the trained shadow target and reference models, leading to performance variance in the proposed attack. To evaluate this factor, we divide an auxiliary dataset $\mathcal{D}_{aux}$ consisting of 25k data records into two parts: $\mathcal{D}^{train}_{shadow}$ for training the shadow target model and $\mathcal{D}^{train}_{ref}$ for training the reference model. We set up two configurations to vary the sizes of $\mathcal{D}^{train}_{shadow}$ and $\mathcal{D}^{train}_{ref}$. In the first configuration, we fix $\mathcal{D}^{train}_{shadow}$ to be $1/5$ of $\mathcal{D}_{aux}$ and vary the size of $\mathcal{D}^{train}_{ref}$. In the second configuration, we do the opposite by fixing $\mathcal{D}^{train}_{ref}$ to be $1/5$ of $\mathcal{D}_{aux}$ and vary the size of $\mathcal{D}^{train}_{shadow}$. We then evaluate the TPR at $1\%$ FPR of LDC-MIA on the VGG-16 target model trained with the CIFAR-10 dataset. The results are shown in Figure 11. Note that the two subfigures have the same label on $y-$axis.
@@
The figure shows that increasing the size of the datasets, $\mathcal{D}^{train}_{shadow}$ and $\mathcal{D}^{train}_{ref}$, improves the TPR of LDC-MIA. When the size of $\mathcal{D}^{train}_{shadow}$ is increased, the MIA classifier is exposed to a wider range of membership scores and more calibrated membership scores from the target samples, which in turn improves the generalization of the shadow target model. On the other hand, when the size of $\mathcal{D}^{train}_{ref}$ is increased, the attacker can obtain more accurate calibrated membership scores by improving the discriminative power of the reference model. The figures also show that the impact of the size of $\mathcal{D}^{train}_{shadow}$ is similar to that of the size of $\mathcal{D}^{train}_{ref}$.
5.4 Model Architectures
Attackers in real-world scenarios may not have knowledge about the architecture of the target model. To analyze the impact of having different network architectures in the shadow target model on the performance of our proposed attack, we randomly select two models from VGG-16, ResNet-18, ResNet-34, and Smallnet as the target and shadow target models while keeping the reference model’s architecture the same as the shadow target model. We evaluate the performance of LDC-MIA using AUC and TPR at $1\%$ FPR as metrics, and we use the CIFAR-10 dataset in this experiment. The results are shown in Figure 12.
@@ 5.4 Model Architectures
Different shapes are used in the figure to represent different shadow target model architectures. It can be observed from the figure that, in general, the architecture of the shadow target model has little impact on the attack performance. For TPR values, we can see that LDC-MIA performs the best when the shadow target model shares the same architecture as the target model from Figure 11(a). However, for AUC, LDC-MIA can achieve the highest value when the two model architectures are different in some cases, as demonstrated in Figure 11(b). For instance, when the target model is VGG-16, the best-performing shadow target model is ResNet-18. It is worth noting that even in the worst cases, LDC-MIA still achieves better TPR than the other MIA methods we have compared with, as shown in Table II. These results are encouraging as they indicate that the attackers are not limited to specific model architectures with our proposed attack.
@@ 5.5 Model Learning Optimizers
There are various choices of optimizers for the training of machine learning models, such as SGD, SGDM, and ADAM. Some optimizers provide better regularization, leading to better generalization and reduced overfitting. For instance, ADAM, a commonly used optimizer, is considered helpful in mitigating memorization. In this experiment, we investigate (1) What impact does the optimizer’s choice have on the attack performance? and (2) Does knowing which optimizer is used in the target model improve attack performance? We use VGG-16 model trained on CIFAR-10 dataset for this experiment.
@@
Figure 13: The impact of different training algorithms.
Figure 13 indicates that there is no significant effect on attack performance by varying optimizers for training the target model. The figure also shows that for every optimizer we test, the attacker can achieve the highest TPR by applying the same optimizer in the shadow target model as in the target model. Applying SGDM in the shadow target model always tends to achieve better attack performance if the exact optimizer in the target model is unknown to the attacker.
@@ 5.6 Different Features
To evaluate the impact of the features introduced into the classifier, we remove each of them from the model and compare the performance of the proposed attack. The proposed classifier has three features — membership scores on the target model, calibrated membership scores, and labels. We compare the contribution of each feature to the attack by analyzing the full log-scale ROC curves. The target model in this experiment is VGG-16 trained on CIFAR-10. The results are shown in Figure 14. From the figure, we can see that removing any of the features results in a degraded performance of the attack, and each of them contributes differently to the attack.
@@
Firstly, removing the label feature leads to a reduction in TPR at low FPR. This indicates that the label feature helps reduce false positives, leading to improved TPR, especially at low FPR. Secondly, removing the membership score reduces TPR in all FPR regions. This verifies what we discussed in Section 3.2 — that including the membership scores not only helps identify hard-to-predict members but also easy-to-predict members, thus improving TPR in all FPR regions. Finally, the performance of the proposed attack significantly degrades by excluding the calibrated membership scores. This is because the calibrated membership scores help separate the hard-to-predict members from the easy-to-predict non-members, which is a significant portion of the non-members, easily. Overall, all three features contribute significantly to the success of our proposed attack.
@@ 7 Conclusion
In this paper, we delve into the difficulty calibration based MIAs and propose a novel learning-based attack, called LDC-MIA. This attack improves the performance of MIA, particularly the TPRs at low FPRs, by using features that characterize the hardness levels of data records. To achieve this, we leverage target samples’ labels, neighborhood information, calibrated membership score, and membership score on the target model. Our experiments show that LDC-MIA can achieve state-of-the-art performance in terms of TPRs at low FPRs, AUC, and precision at high recall rates while keeping the attack cost low.

### v3

@@ 4.5 Effect of data augmentation
Data augmentation is a technique that can be used to increase the size of a dataset by applying different transformations to existing training data. This technique is often used to improve the generalization and robustness of models. By exposing the model to a wider range of data variations, it can learn to handle different input scenarios, resulting in better performance. In this experiment, we applied data augmentation during the training of the target model and the shadow models (if any). Then, we evaluated its effect on the MIAs by measuring its AUC and TPR at $1\%$ FPR values. Since attackers may not have access to the data augmentation techniques used in the target model in real-life situations, we train the shadow target model with random data augmentations. Specifically, we use horizontal flipping for the shadow (target) model training and random cropping and rotation for the target model training. When querying the target model, we use the original target sample. The experiments are conducted on the CIFAR-10 and CIFAR-100 datasets. The target model for CIFAR-10 is VGG-16, and the target model for CIFAR-100 is SmallNet. The experiment results are shown in Table VII.
TABLE VII: The impact of data augmentation
@@
w/o aug w/ aug w/o aug w/ aug
AUC TPR AUC TPR AUC TPR AUC TPR
Salem et al. [29] 0.679 1.2% 0.605 0.2% 0.697 2.3% 0.625 0.6%
Yeom et al. [39] 0.696 0.0% 0.602 0.0% 0.889 1.0% 0.802 0.0%
Watson et al. [37] 0.705 3.4% 0.651 1.1% 0.772 5.9% 0.709 2.7%
Ours 0.752 6.7% 0.712 3.8% 0.918 16.4% 0.869 12.4%
@@ 5.1 Differential Privacy
In order to evaluate the robustness of our proposed attack, we utilize the concept of differential privacy during the training of the target model. This technique adds noise or randomization to the data, which helps protect individual privacy in datasets. It can be useful in limiting the effectiveness of many existing MIAs [37, 3, 39]. We use DP-SGD [1], one of the state-of-the-art DP mechanisms, for training the target model in our experiments. DP-SGD adds carefully calibrated noise to the gradients computed during each iteration. The amount of noise added depends on the sensitivity of the gradients and the desired privacy budget $\varepsilon$. While a smaller $\varepsilon$ provides stronger privacy guarantees, it can also result in noisier updates. The other two important parameters in DP-SGD are the clipping bound $C$ and the noise multiplier $\sigma$.
@@
Figure 10 and Table VIII show that as the desired privacy budget $\varepsilon$ increases, both AUC and TPR decrease. However, in practice, there is a trade-off between privacy ($\varepsilon$) and utility (the accuracy of the trained model). When higher privacy is required, adding more noise can significantly affect the model’s utility. This is evident from the model accuracy column in Table VIII. In practice, to preserve model accuracy, reasonable values of $\varepsilon$, such as 100 or 1000, are more likely to be used. It is observed that AUC and TPR are not significantly reduced in these cases. Moreover, even with small $\varepsilon$ values, the proposed attack can still achieve high precision values ($>90\%$) at low recall, as seen from Figure 10(b).
@@ 5.2 Overfitting Level of the Target Model
Previous studies [29, 31] have demonstrated that the performance of MIAs is closely related to the overfitting level of the target model. Overfitting occurs when the model fits the training data too well, even with noise and unique patterns to its training dataset. Several factors can affect a model’s overfitting level, such as dataset size and quality, training rounds, model complexity, and regularization. Typically, increasing the training dataset’s size helps reduce overfitting, while decreasing it has the opposite effect. Because a larger and more diverse dataset allows the model to observe a broader range of variations and generalize better with less memorization of specific data points.
In our experiments, we adjust the training dataset size between 6500 and 12500 to vary the target model’s overfitting level. Meanwhile, we keep the size of the training dataset of the reference model $\mathcal{D}^{train}_{ref}$ and that of the shadow target model $\mathcal{D}^{train}_{shadow}$ fixed. We then measure the impact of the overfitting level by evaluating AUC and TPR of the proposed attack. The results are shown in Table IX.
@@
Table IX depicts that the overfitting level of the target model increases as the size of its training dataset decreases, by looking at the gap between the training accuracy and test accuracy of the target model. The larger the gap between these two, the more the model is overfitting. We note that in our experiment for our method LDC-MIA the AUC and TPR at $1\%$ FPR of the MIA improve slightly as the overfitting level of the target model increases. Membership inference attack benefits from a higher level of overfitting, which could mean a higher level of memorization. The above results align with the findings on other MIAs in the literature.
@@ 5.3 Training Dataset Sizes for the Shadow Target and the Reference Models
The size of the datasets used to train the shadow target and reference models is an important factor in our proposed attack. As discussed in Section 5.2, the training dataset sizes affect the performance of the trained shadow target and reference models, leading to performance variance in the proposed attack. To evaluate this factor, we divide an auxiliary dataset $\mathcal{D}_{aux}$ consisting of 25k data records into two parts: $\mathcal{D}^{train}_{shadow}$ for training the shadow target model and $\mathcal{D}^{train}_{ref}$ for training the reference model. We set up two configurations to vary the sizes of $\mathcal{D}^{train}_{shadow}$ and $\mathcal{D}^{train}_{ref}$. In the first configuration, we fix $\mathcal{D}^{train}_{shadow}$ to be $1/5$ of $\mathcal{D}_{aux}$ and vary the size of $\mathcal{D}^{train}_{ref}$. In the second configuration, we do the opposite by fixing $\mathcal{D}^{train}_{ref}$ to be $1/5$ of $\mathcal{D}_{aux}$ and vary the size of $\mathcal{D}^{train}_{shadow}$. We then evaluate the TPR at $1\%$ FPR of LDC-MIA on the VGG-16 target model trained with the CIFAR-10 dataset. The results are shown in Figure 11. Note that the two subfigures have the same label on $y-$axis.
@@
Figure 13: The impact of different training algorithms.
Figure  11 shows that increasing the size of the datasets, $\mathcal{D}^{train}_{shadow}$ and $\mathcal{D}^{train}_{ref}$, improves the TPR of LDC-MIA. Increased size of $\mathcal{D}^{train}_{shadow}$ improves the generalization of the shadow target model. The MIA classifier benefits from the better shadow target model due to exposure to a broader range of membership scores and calibrated membership scores. On the other hand, the increased size of $\mathcal{D}^{train}_{ref}$ improves the reference model, which leads to the improvement of the MIA classifier through more accurate calibrated membership scores. Figure 11 also shows that the effect on the MIA classifier performance of the size of $\mathcal{D}^{train}_{shadow}$ is similar to that of the size of $\mathcal{D}^{train}_{ref}$.
5.4 Model Architectures
Let’s consider a real-world threat scenario where the attackers do NOT have knowledge about the architecture of the target model, so they guess an architecture for auxiliary models. We want to analyze the impact of having different network architectures in the shadow target model on the performance of LDC-MIA. To do so we randomly select two models from VGG-16, ResNet-18, ResNet-34, and Smallnet as the target and shadow target models while keeping the reference model’s architecture the same as the shadow target model. We evaluate the performance of LDC-MIA using AUC and TPR at $1\%$ FPR as metrics, on the CIFAR-10 dataset. The results are shown in Figure 13.
Figure 13 suggests that the architecture of the shadow target model does NOT have much impact on the attack performance. TPR values of LDC-MIA are the best when the shadow target model shares the same architecture as the target model. However, AUC of LDC-MIA can achieve the highest value when the two model architectures differ in some cases, as depicted in Figure 12(b). For instance, when the target model is VGG-16, the best-performing shadow target model uses ResNet-18. These results suggest that for  LDC-MIA there is no absolute need to know the specific model architectures of the target model to launch an equivalently successful attack. Although we note that we use a limited number of pre-defined candidate architectures for guessing in this experiment, therefore the results are indicative but not comprehensive.
@@ 5.5 Model Learning Optimizers
There are more than one optimizers for training the machine learning models, Some widely-use examples include SGD, SGDM, and ADAM. Some optimizers provide better regularization, leading to better generalization and reduced overfitting. For instance, ADAM, a commonly used optimizer, is considered helpful in mitigating memorization. We use VGG-16 model on CIFAR-10 dataset in an experiment to investigate (1) What impact does the optimizer have on the attack performance? and (2) Does knowing which optimizer is used in the target model improve attack performance?
@@
Figure 13 indicates that there is no significant effect on attack performance by varying optimizers for training the target model. The figure also shows that for every optimizer we test, the attacker can achieve the highest TPR by applying the same optimizer in the shadow target model as in the target model. Interestingly, in our experiment applying SGDM in the shadow target model consistently achieves better attack performance if the exact optimizer in the target model is unknown to the attacker.
@@ 5.6 Different Features
To evaluate the impact of the features introduced into the classifier in LDC-MIA, we remove one feature each time from the model training and compare the performance of the resulted classifiers that are trained on all-but-one features. The MIA classifier in LDC-MIA has three features — membership scores on the target model, calibrated membership scores, and labels. We compare the contribution of each feature to the attack by analyzing the full log-scale ROC curves. The target model in this experiment is VGG-16 trained on CIFAR-10. The results are shown in Figure 14. The figure indicates that removing any of the features results in degraded attack performance, and each feature contributes to the attack in different ways.
@@
Firstly, removing the label feature leads to a reduction in TPR at low FPR. This indicates that the label feature helps reduce false positives, leading to improved TPR, especially at low FPR. Secondly, removing the membership score reduces TPR in all FPR regions. This verifies what we discussed in Section 3.2 — that including the membership scores not only helps identify hard-to-predict members but also easy-to-predict members, thus improving TPR in all FPR regions. Finally, the performance of the proposed attack significantly degrades by excluding the calibrated membership scores. This is because the calibrated membership scores help separate the hard-to-predict members from the easy-to-predict non-members, which is a significant portion of the non-members, easily. Overall, all three features offer unique contributions to the success of our proposed attack.
@@ 7 Conclusion
In this paper, we delve into the difficulty calibration based MIAs and propose a novel learning-based attack, called LDC-MIA. This attack improves the performance of MIA, particularly the TPRs at low FPRs, by using features that characterize the hardness levels of data records. To achieve this, we leverage target samples’ labels, neighborhood information, calibrated membership score, and membership score on the target model. Our experiments show that LDC-MIA can achieve state-of-the-art performance in terms of TPRs at low FPRs, AUC, and precision at high recall rates while keeping the attack cost relatively low.
8 Acknowledgement
We would like to thank our sheperd and the reviewers of Euro S&P’24 for their invaluable feedback. This work is partially supported by an NSF grant CNS-2008468 and an ONR grant N00014-23-1-2137.